i^103
how do I simplify this?
something about dividing by four and taking the remainder??

Answer by user_dude2008(1862)

i^103 = i^(100+3)=i^100*i^3 = (i^4)^25*i^3 = (1)^25*i^3 = i^3 = -i

Answer: i^103 = -i